OpenStudy (anonymous):

charmaine drove 715 miles in 11 hours. at the same rate, how long would it take her to drive 845 miles? don't understand

OpenStudy (bibby):

d=r*t find the rate

OpenStudy (anonymous):

oh so 715 / 11 = 65 845 / 65 = 13

TheSmartOne (thesmartone):

Yes.

TheSmartOne (thesmartone):

Good job @WeMaejor900

OpenStudy (anonymous):

thanks & hey does anyone know what math for college readiness is?

OpenStudy (bibby):

didn't have that when I was in high school

TheSmartOne (thesmartone):

nope..

OpenStudy (anonymous):

im not sure, but i think this class is based on where you live. what math classes you took in high school

OpenStudy (anonymous):

@bibby im not sure, but i think this class is based on where you live. what math classes you took in high school

OpenStudy (bibby):

what about it?

OpenStudy (anonymous):

@bibby because I think based off where someone lives they might have different math credits or something.

OpenStudy (bibby):

math is math no matter what the subject is called

OpenStudy (anonymous):

@bibby true but thanks anyway